OpenStudy (solomonzelman):

I didn't get any help all the previous times, can s/o help me at this at least? Prove that sin^2x+cos^2x=1 Please don't give me any links; explain!!!!

OpenStudy (amistre64):



OpenStudy (amistre64):

by definitions: sinx = b/r; b = r sinx cosx = a/r; a = r cosx by pythag thrm a^2 + b^2 = r^2 substitute and simplify

OpenStudy (perl):

@amistre64 so let r = 1 ?

OpenStudy (amistre64):

letting r=1 is not general enough

OpenStudy (solomonzelman):

Guys guys I am trying to get SAT practice questions.... PLEASE !!! I was like dude, learn how to read please...(Sorry I had to say that....

OpenStudy (perl):

@amistre64 ok i see what you did, very nice argument. you can also do the following. let x = theta, then use definition of sin theta and cos theta using a unit cirlce

OpenStudy (nincompoop):

it's the same

OpenStudy (perl):

right, its similar. draw a circle with radius r

OpenStudy (solomonzelman):

what?! All right I am out I have to ace the English test and come back here Good buy!!

OpenStudy (nincompoop):

that's about the most straight-forward proof, @SolomonZelman

OpenStudy (perl):

you can draw a circle with radius 1, that will work too

OpenStudy (perl):